### v1

@@ 3 Illustration of Infeasibility Phenomenon
Figure 2: State trajectories and state feasibility of MPC under pointwise constraints.
4 Feasibility and Feasible Regions
We have seen that the improper selection of virtual-time constraints would result in the notorious infeasibility phenomenon. Now let us move to the formal definition of feasibility, which allows us to analyze what kind of virtual-time constraint can guarantee safety. The aforementioned example of pointwise constraint enlightens us that depending on the choice of constraint, virtual-time OCPs may suddenly lose feasibility at some time points or keep feasibility forever. These two results bring about the need to distinguish between (1) initial feasibility and (2) endless feasibility. As suggested by their names, initial feasibility may be only temporary, while endless feasibility always comes with all virtual-time OCPs in the everlasting future.
@@ 4.1 Initial feasibility
The existence of infeasibility phenomenon indicates that a state with feasible virtual-time OCP may inevitably evolve into a state with infeasible virtual-time OCP. As a result, the property of a state’s yielding feasible OCP at the current step is termed as initial feasibility. This property is just a shortsighted feature and may not give any guarantee in the future (even just one step later!). To have an intuitive understanding of initial feasibility, recall the evolution of infeasibility phenomenon in Figure 1, where $x_{t}$ is a feasible state but still evolves into a state $x_{t+2}$ that is infeasible. The definition of initial feasibility is as follows.
Definition 4.1 (Initial feasibility of a state).
@@
The subscript “init” of initial feasible region $\mathrm{X}_{\mathrm{init}}^{g}$ stands for “initial”, and the superscript “$g$” emphasizes that it is related to the virtual-time constraint. For a state to be initially feasible, the only requirement is that the virtual-time OCP starting from it has a solution. Outside the IFR, there is no policy satisfying the virtual-time constraint, and hence the virtual-time OCP has no solution. In other words, IFR is a region restricted by the chosen virtual-time constraint, identifying those states that are meaningful under the virtual-time constraint. In this definition, the policy $\pi$ can be arbitrarily chosen to meet the existence statement. Supposing that we have already been given a policy, it is also a natural need to check whether its output satisfies the virtual-time constraint, which leads to the definition of initial feasibility of a policy.
Definition 4.2 (Initial feasibility of a policy).
@@
The notation $\mathrm{X}_{\mathrm{init}}^{g}(\pi)$ means that the IFR is not only dependent on the constraint $g$ but also a function of policy $\pi$. The initial feasibility of a policy is closely related to the initial feasibility of a state. If a state is initially feasible, then there must exist a policy that is initially feasible in this state. On the other hand, if a policy is initially feasible in a state, then there exists at least one solution to virtual-time OCP starting from this state, which indicates that this state is initially feasible. The IFR of a policy $\mathrm{X}_{\mathrm{init}}^{g}(\pi)$ is the region where one is approved to take policy $\pi$ under the virtual-time constraint. Similar to the initial feasibility of a state, the initial feasibility of a policy only describes the property of a single virtual-time OCP starting from the current real-time step. It does not provide any guarantee on the virtual-time OCPs in future real-time steps. A policy that is initially feasible at the current time step may become infeasible after it takes a step forward in the real-time domain. Examples of IFR of a policy are the union of blue and red regions in Figure 1 and Figure 2. Within these regions, the virtual-time constraints are satisfied by the policy, but there is no guarantee that they will still be satisfied in successive states.
@@ 4.2 Endless feasibility
Having the definition of initial feasibility at hand, we are able to describe what we actually care about, i.e., a long-term property taking infinite horizon constraints into account. This property is termed as endless feasibility. Intuitively, endless feasibility gives a guarantee of not only the feasibility of the current virtual-time OCP but also that of all future virtual-time OCPs. Here is an example to give a sense of why there is a distinction between initial and endless feasibility. In a car-following scenario, the ego vehicle is required to follow a leading vehicle while keeping a safe distance from it. Suppose that the virtual-time constraint requires a positive distance between vehicles only at the next time step. Then, a state with an overly high speed may be initially feasible since the vehicles will not collide immediately. However, due to the limited braking capability, the car will inevitably reach a state where collision is doomed at the next time step no matter what action is taken. In other words, the virtual-time OCP becomes infeasible after certain steps. Unlike initial feasibility, endless feasibility guarantees the feasibility of all future virtual-time OCPs by its definition. Just like the case for initial feasibility, it can also be defined for a state or a policy.
Definition 4.3 (Endless feasibility of a state).
@@
The subscript “edls” in the above symbols is an abbreviation of “endless”, and the superscript “$g$” emphasizes that it is related to the virtual-time constraint. The endless feasibility of a state is related to not only the current step in the real-time domain but also all future time steps. The states in future time steps can be obtained by arbitrary initially feasible policies in each time step. The arbitrariness of policy selection is critical to the definition of endless feasibility. This is because, for any initially feasible state, there may exist more than one initially feasible policy in that state. Which one is the solution of the virtual-time OCP depends on the objective function. The arbitrariness of policy selection ensures that the solution can always ensure the successive states to be initially feasible, regardless of the objective function. As a result, the EFR identifies those states that can be definitely rendered safe by the virtual-time constraint. Similar to the initial feasibility of a policy, we can also define the endless feasibility of a policy.
Definition 4.4 (Endless feasibility of a policy).
@@
The “$\pi$” in brackets serves the same purpose as in $\mathrm{X}_{\mathrm{init}}^{g}(\pi)$. Note that the definition of endless feasibility is built upon initial feasibility, requiring all successive states in a trajectory to be initially feasible. For endless feasibility of a policy, this trajectory is naturally induced by the given policy. For endless feasibility of a state, this trajectory can be induced by any initial feasible policy, i.e., it is required that at every step, taking any initially feasible policy (not just one of them), the next state is also initially feasible. With such a definition, it naturally holds that for an endlessly feasible state, there must also be an endlessly feasible policy in that state, i.e., EFR is a subset of the policy’s EFR. This conclusion will be formally stated and proved in Theorem 5.1. Examples of EFR of a policy are the red regions in Figure 1 and Figure 2. Within these regions, the policy is initially feasible in both the current and all successive states.
@@
The above-mentioned four definitions are all specific to virtual-time constraints. Even for the same real-time OCP, a different choice of virtual-time constraint will lead to different feasible regions. A natural question is how to design a virtual-time constraint that induces an EFR as large as possible. To discuss this, we first need to define the maximum EFR.
Definition 4.5.
@@
The maximum endlessly feasible region (shortened as maximum EFR), denoted as $\mathrm{X}_{\mathrm{edls}}^{*}$, is the union of all endlessly feasible regions.
@@
The above definition of maximum EFR is closely related to the notion of maximum control invariant set in MPC. The latter refers to the maximum set in which for any state, there exists a control sequence that keeps its successive states still in this set. The maximum EFR has one more requirement than the maximum control invariant set, i.e., all states in it must satisfy the real-time constraint. In other words, a constraint-satisfying maximum control invariant set is a maximum EFR. Either too strong or too weak a virtual-time constraint will induce an EFR $\mathrm{X}_{\mathrm{edls}}^{g}$ smaller than $\mathrm{X}_{\mathrm{edls}}^{*}$. For the former case, less state is considered initially feasible and hence outside of $\mathrm{X}_{\mathrm{edls}}^{g}$. For the latter case, one may take too aggressive actions at the beginning of a trajectory due to the inadequate restriction of virtual-time constraint, leading to the occurrence of an infeasibility phenomenon. Examples of these two cases can be found in the experiment results in Section 8, where the pointwise constraint with $n=2$ is too weak, and the control barrier function constraint with $k=0.5$ is too strong. The maximum EFR is the largest area in which we can expect any policy to work safely. The following theorem gives an identical description of the maximum EFR, which explicitly relates endless feasibility with constraint satisfaction in real-time domain.
Theorem 4.1.
@@
Proof.
First, we prove 1) $\Rightarrow$ 2). According to Definition 4.5,
@@
as the virtual-time constraint. In this case, we have $x\in\mathrm{X}_{\mathrm{edls}}^{g}$. Since $\mathrm{X}_{\mathrm{edls}}^{g}\subseteq\mathrm{X}_{\mathrm{edls}}^{*}$, we conclude that $x\in\mathrm{X}_{\mathrm{edls}}^{*}$. ∎
The above theorem justifies the importance of defining maximum EFR in that states in $\mathrm{X}_{\mathrm{edls}}^{*}$ are exactly what we are eager to identify, which are the only states that can be rendered safe in the real-time domain. That is to say, for any state $x\in\mathrm{X}_{\mathrm{edls}}^{*}$, there exists a policy $\pi$, such that all future states are in the constrained set $\mathrm{X}_{\mathrm{cstr}}$. Since 2) in Theorem 4.1 goes to infinity, the successive states are always in the maximum EFR.
Intuitively, if a state $x$ is in the EFR of a policy $\pi$, it can be rendered safe by this policy. This result is stated formally in the following theorem, which theoretically justifies applying $\pi$ in $x$ and all successive states.
Theorem 4.2.
@@
Proof.
According to Definition 4.4,
@@
where $x_{t+i}$ are obtained by $\pi$. Since $\mathrm{X}_{\mathrm{init}}^{g}\subseteq\mathrm{X}_{\mathrm{cstr}}$, we have $x_{t+i}\in\mathrm{X}_{\mathrm{cstr}}$, i.e., $h(x_{t+i})\leq 0$. ∎
Theorem 4.2 tells us that the EFR of a policy is where we can safely apply the policy in real-time domain. This is the reason why we need to find not only the policy itself but also its EFR when solving constrained OCPs. In literature, this EFR is found by identifying a control certificate related to virtual-time constraint, e.g., control barrier function (Ames et al., 2014), safety index (Liu and Tomizuka, 2014), and Hamilton-Jacobi reachability function (Mitchell et al., 2005). In fact, initial feasibility is not what we ultimately care about because it only describes whether constraints of the current virtual-time OCP are satisfied. The role of initial feasibility is to help us define endless feasibility, which describes whether constraints of all successive virtual-time OCPs are satisfied. Our main concern is endless feasibility because it is related to long-term safety in real-time domain. Therefore, we care about how to calculate the EFR of a policy, especially the policy obtained from solving a virtual-time OCP. We hope that the EFR of this policy equals the maximum EFR. Furthermore, the size of EFR of a policy is related to its performance. If we want a policy to be optimal, then its EFR must contain all optimal state trajectories. A larger EFR has a greater chance of containing optimal trajectories and therefore increases the probability of finding the optimal policy.
5 Some Properties of Feasible Regions
In this section, we formally explore the properties of feasible regions, as well as their relationships and how they are informative for virtual-time constraint design. Specifically, we will pay attention to containment relationships and equivalence relationships of different feasible regions.
@@ 5.1 Containment relationships of feasible regions
The definitions of feasible regions can be categorized by two dimensions: (a) initial ones or endless ones and (b) with a given policy or without a given policy. The four combinations correspond to (1) initially feasible region (IFR) $\mathrm{X}_{\mathrm{init}}^{g}$, (2) IFR of a policy $\mathrm{X}_{\mathrm{init}}^{g}(\pi)$, (3) endlessly feasible region (EFR) $\mathrm{X}_{\mathrm{edls}}^{g}$, and (4) EFR of a policy $\mathrm{X}_{\mathrm{edls}}^{g}(\pi)$. Besides these, there are two more regions we need to pay attention to, i.e., constrained set $\mathrm{X}_{\mathrm{cstr}}$ and the maximum EFR $\mathrm{X}_{\mathrm{edls}}^{*}$. The constrained set comes from real-world constraints, and the maximum EFR is the feasible region that we always aim to identify because it gives the largest allowable working area of the policy. Among all of these six regions, $\mathrm{X}_{\mathrm{cstr}}$, $\mathrm{X}_{\mathrm{init}}^{g}$, and $\mathrm{X}_{\mathrm{init}}^{g}(\pi)$ are related only to the current step, while $\mathrm{X}_{\mathrm{edls}}^{g}$, $\mathrm{X}_{\mathrm{edls}}^{g}(\pi)$, and $\mathrm{X}_{\mathrm{edls}}^{*}$ are our true ultimate goals. Since the latter is related to long-term feasibility, their identification is much more difficult than that of the former. The following theorem gives some containment relationships between these six regions so that we can let those easy-to-identify regions cast some light on the goal regions.
Theorem 5.1 (Feasible region containment).
@@
    The first containment relationship $\mathrm{X}_{\mathrm{edls}}^{g}\subseteq\mathrm{X}_{\mathrm{init}}^{g}$ can be directly concluded from Definition 4.3. For the second one, according to Definition 4.1,
@@
    This holds by Definition 4.4 and Definition 4.2.
@@
    where $x_{t}=x$. Thus, $x_{t+i}\in\mathrm{X}_{\mathrm{init}}^{g}(\pi)$. Since $\mathrm{X}_{\mathrm{init}}^{g}(\pi)\subseteq\mathrm{X}_{\mathrm{init}}^{g}$, it follows that $x_{t+i}\in\mathrm{X}_{\mathrm{init}}^{g}$. Thus, $x_{t+i}\in\mathrm{X}_{\mathrm{cstr}}$, i.e., $h(x_{t+i})\leq 0$. Recall Theorem 4.1 and we reach $\mathrm{X}_{\mathrm{edls}}^{g}(\pi)\subseteq\mathrm{X}_{\mathrm{edls}}^{*}$.
@@
    Since $\mathrm{X}_{\mathrm{init}}^{g}=\mathrm{X}_{\mathrm{init}}^{g}(\pi)$, it holds that $\pi$ is initially feasible in all states in $\mathrm{X}_{\mathrm{init}}^{g}$. That is to say, $\pi\in\Pi_{\mathrm{init}}^{g}(x)$,$\forall x\in\mathrm{X}_{\mathrm{init}}^{g}$. Starting from any $x\in\mathrm{X}_{\mathrm{edls}}^{g}$ and choosing
@@
Figure 3 illustrates the containment relationships in Theorem 5.1. Figure 3(a) is an illustration of $\mathrm{X}_{\mathrm{edls}}^{g}\subseteq\mathrm{X}_{\mathrm{init}}^{g}\subseteq\mathrm{X}_{\mathrm{cstr}}$. The light-colored circles correspond to virtual-time states, and the dark-colored ones stand for real-time states. The first containment relationship holds by definition since a state must first be initially feasible, then it could be endlessly feasible. This is also the case for the second one.
Figure 3(b) is an illustration of $\mathrm{X}_{\mathrm{init}}^{g}(\pi)\subseteq\mathrm{X}_{\mathrm{init}}^{g}$. For a state in $\mathrm{X}_{\mathrm{init}}^{g}$, there exists a policy satisfying the virtual-time constraint, but it is not necessarily the given policy $\pi$. Intuitively, given a policy $\pi$ poses a greater restriction on the IFR, making the IFR of a policy become a subset of IFR.
Figure 3(c) is an illustration of $\mathrm{X}_{\mathrm{edls}}^{g}(\pi)\subseteq\mathrm{X}_{\mathrm{edls}}^{*}$. On the left are the zero-sublevel set of $h(x)$ and $g(x)$, they correspond to the maximum EFR $\mathrm{X}_{\mathrm{edls}}^{*}$ and the EFR of a policy $\mathrm{X}_{\mathrm{edls}}^{g}(\pi)$ on the right, respectively. To see how $\{x|h(x)\leq 0\}$ is related to $\mathrm{X}_{\mathrm{edls}}^{*}$, recall Theorem 4.1 which states that $x\in\mathrm{X}_{\mathrm{edls}}^{*}$ is equivalent to the existence of a policy rendering $x$ safe in real-time domain. Likewise, $x\in\mathrm{X}_{\mathrm{edls}}^{g}(\pi)$ means that the specific policy $\pi$ renders $x$ safe in virtual-time domain, without violating virtual-time constraint $g(x)\leq 0$. Naturally, a larger valid state set ($\{x|h(x)\leq 0\}$ versus $\left\{x|g(x)\leq 0\right\}$) and no restriction on the policy yields a larger EFR.
@@
Figure 3(d) is an illustration of the statement “If $\mathrm{X}_{\mathrm{init}}^{g}(\pi)=\mathrm{X}_{\mathrm{init}}^{g}$, then $\mathrm{X}_{\mathrm{edls}}^{g}\subseteq\mathrm{X}_{\mathrm{edls}}^{g}(\pi)$.” The left picture corresponds to the condition $\mathrm{X}_{\mathrm{init}}^{g}(\pi)=\mathrm{X}_{\mathrm{init}}^{g}$, with the red region representing both $\mathrm{X}_{\mathrm{init}}^{g}$ and $\mathrm{X}_{\mathrm{init}}^{g}(\pi)$. This condition implies that $\pi\in\Pi_{\mathrm{init}}^{g}(x),\forall x\in\mathrm{X}_{\mathrm{init}}^{g}$, where $\Pi_{\mathrm{init}}^{g}(x)$ is the set of policies that are initially feasible under $x$. The right picture illustrates how $\mathrm{X}_{\mathrm{edls}}^{g}$ is a subset of $\mathrm{X}_{\mathrm{edls}}^{g}(\pi)$. For a state in the purple region $\mathrm{X}_{\mathrm{edls}}^{g}$, every initially feasible policy at every time step must lead to an initially feasible state at the next time step. That is to say, the policy $\pi^{\prime}$ (in purple) and $\pi^{\prime\prime}$ (in red) are chosen arbitrarily from the initially feasible policy sets $\Pi_{\mathrm{init}}^{g}(x_{t+1})$ and $\Pi_{\mathrm{init}}^{g}(x_{t+2})$, respectively. So $\pi$ is also a possible choice for $\pi^{\prime}$ and $\pi^{\prime\prime}$ (and all future policies). This guarantees that $\pi$ renders any state in $\mathrm{X}_{\mathrm{edls}}^{g}$ endlessly feasible under itself.
Figure 3: Illustration of containment relationships.

### v2

@@ 2.2 Real-time domain and virtual-time domain
Figure 3.1: State trajectories and state feasibility of MPC under pointwise constraints. The circles stand for the states at every step on the trajectories, which are obtained by solving problem (2.5) in an receding-horizon manner. The squares and diamonds stand for regular states so as to show different feasibility regions. Regardless of the shape, all red marks represent endlessly feasible states, all blue marks for initially feasible states, and all gray marks for infeasible states. The purple-shaded areas represent real-time constraint.
Chapter 4 Feasibility and Feasible Regions
We have seen that improper design of virtual-time constraints will result in the infeasibility phenomenon. Now, let us move to the formal definition of feasibility, which allows us to analyze what kind of virtual-time constraint can guarantee safety. The aforementioned example of pointwise constraint suggests us that depending on the choice of constraint, virtual-time OCPs may suddenly lose feasibility at some time points or keep feasibility forever. These two results bring about the need to distinguish between 1) initial feasibility and 2) endless feasibility. As suggested by their names, initial feasibility is only temporary, while endless feasibility regards all virtual-time OCPs in the everlasting future.
@@ 4.1 Initial feasibility
The property that a state yields a feasible OCP at the current step is termed as initial feasibility. This property is just a shortsighted feature and may not give any guarantee in the future (even just one step later!). To understand this, recall the evolution of infeasibility phenomenon in Figure 1.1, where $x_{t}$ is a feasible state but still evolves into an infeasible state $x_{t+2}$. The definition of initial feasibility is as follows.
Definition 4.1.1 (Initial feasibility of a state).
@@
The subscript “init” of initial feasible region $\mathrm{X}_{\mathrm{init}}^{g}$ stands for “initial”, and the superscript “gg” emphasizes that it is related to the virtual-time constraint. For a state to be initially feasible, the only requirement is that the virtual-time OCP starting from it has a solution. In this definition, the policy $\pi$ can be arbitrarily chosen to meet the existence statement. Supposing that we have already been given a policy, it is also a natural need to check whether the policy satisfies the virtual-time constraint, which leads to the definition of initial feasibility of a policy.
Definition 4.1.2 (Initial feasibility of a policy).
@@
The notation $\mathrm{X}_{\mathrm{init}}^{g}(\pi)$ means that this region is a function of policy $\pi$. The initial feasibility of a policy is closely related to the initial feasibility of a state. If a state is initially feasible, then there must exist a policy that is initially feasible in this state. On the other hand, if a policy is initially feasible in a state, then there exists at least one solution to virtual-time OCP starting from this state, which indicates that this state is initially feasible. Examples of IFR of a policy are the union of blue and red regions in Figure 1.1 and Figure 3.1. Within these regions, the virtual-time constraints are satisfied by the policy, but there is no guarantee that they will still be satisfied in successive states.
@@ 4.2 Endless feasibility
With initial feasibility, we are able to describe what we actually care about, i.e., a long-term property taking infinite horizon constraints into account. This property is termed as endless feasibility. Intuitively, endless feasibility guarantees that all future virtual-time OCPs are feasible. Here is an example of the distinction between initial and endless feasibility. In a car-following scenario, the ego vehicle is required to follow a leading vehicle while keeping a safe distance from it. Suppose that the virtual-time constraint requires a positive distance between vehicles only at the next time step. Then, a state with an overly high speed may be initially feasible since the vehicles will not collide immediately. However, due to the limited braking capability, the car will inevitably reach a state where collision is doomed at the next time step no matter what action is taken. At this time, the virtual-time OCP becomes infeasible. Like initial feasibility, endless feasibility is also separately defined for a state and a policy.
Definition 4.2.1 (Endless feasibility of a state).
@@
The subscript “edls” in the above symbols is an abbreviation of “endless”. The arbitrariness of policy selection is critical to the definition of endless feasibility. This is because in any initially feasible state, there may exist more than one initially feasible policy. Which one is the solution to the virtual-time OCP depends on the objective function. The arbitrariness of policy selection ensures that the solution can always ensure the successive states to be initially feasible, regardless of the objective function.
Definition 4.2.2 (Endless feasibility of a policy).
@@
Note that the definition of endless feasibility is built upon initial feasibility, requiring all successive states in a trajectory to be initially feasible. For endless feasibility of a policy, this trajectory is naturally induced by the given policy. For endless feasibility of a state, this trajectory can be induced by any initial feasible policy. With such a definition, it naturally holds that for an endlessly feasible state, there must also be an endlessly feasible policy in that state, i.e., EFR is a subset of any policy’s EFR. This conclusion will be formally stated and proved in Theorem 5.1.1. Examples of EFR of a policy are the red regions in Figure 1.1 and Figure 3.1. Within these regions, the policy is initially feasible in both the current and all successive states.
Readers who are familiar with MPC may find the concept of endless feasibility similar to the persistent feasibility in MPC. A persistently feasible set is a set that is forward invariant with respect to the MPC feedback law $\pi_{\text{MPC}}$ [boccia2014stability]. The relation between these two concepts lies in that, given a constrained OCP in MPC, the endlessly feasible region of a $\pi_{\text{MPC}}$, i.e., $\mathrm{X}_{\mathrm{edls}}^{g}(\pi_{\text{MPC}})$, equals the maximum persistently feasible set. This is because $\mathrm{X}_{\mathrm{edls}}^{g}(\pi_{\text{MPC}})$ is forward invariant under $\pi_{\text{MPC}}$, and for any $x$ in the maximum persistently feasible set, $\pi_{\text{MPC}}$ must be initially feasible on $x$ and all the subsequent states. However, this feedback law is a specific policy that is made up with the solutions of a series of OCPs, which makes it closely tied to the constrained OCP. Hence, traditional feasibility analysis pertains to the problem itself, not to the state or policy in isolation, which is inadequate in the RL cases. Note the subtle distinction: persistent feasibility in MPC is typically an existential property, i.e., there exists a control action or policy that keeps successive states initially feasible, whereas our notion of endless feasibility is stronger and is achieved only when the conditions hold for all initially feasible policies. The proposed endless feasibility (and the whole feasibility theory) therefore offers a more fine-grained discussion of feasibility at both state and policy levels. At the state level, the definition of feasibility is independent of any specific policy. At the policy level, the policy under discussion is not constrained to be the optimal one or the (suboptimal) solution of an OCP.
@@
Different choices of virtual-time constraints will lead to different feasible regions. A natural question is how to design a virtual-time constraint that induces an EFR as large as possible. To discuss this, we first need to define the maximum EFR.
Definition 4.2.3.
@@
The maximum endlessly feasible region (shortened as maximum EFR), denoted as $\mathrm{X}_{\mathrm{edls}}^{*}$, is the union of EFRs under all possible virtual-time constraints.
@@
Maximum EFR is closely related to the notion of maximum control invariant set in MPC. The latter refers to the maximum set in which for any state, there exists a control sequence that keeps its successive states still in this set. The maximum EFR further requires that all states in it must satisfy the real-time constraint. Either too strong or too weak a virtual-time constraint will induce an EFR $\mathrm{X}_{\mathrm{edls}}^{g}$ smaller than $\mathrm{X}_{\mathrm{edls}}^{*}$. For the former case, less state is considered initially feasible and hence outside of $\mathrm{X}_{\mathrm{edls}}^{g}$. For the latter case, one may take too aggressive actions at the beginning due to inadequate restriction of virtual-time constraint, leading to the infeasibility phenomenon. Examples of these two cases can be found in Chapter 8, where a pointwise constraint with $n=2$ is too weak, and a control barrier function constraint with $k=0.5$ is too strong. The following theorem provides an equivalent characterization of the maximum EFR (Definition 4.2.3), explicitly relating endless feasibility to constraint satisfaction in the real-time domain.
Theorem 4.2.4.
@@
Proof 4.2.5.
First, we prove 1) $\Rightarrow$ 2). According to Definition 4.2.3,
@@
as the virtual-time constraint. In this case, we have $x\in\mathrm{X}_{\mathrm{edls}}^{g}$. Since $\mathrm{X}_{\mathrm{edls}}^{g}\subseteq\mathrm{X}_{\mathrm{edls}}^{*}$, we conclude that $x\in\mathrm{X}_{\mathrm{edls}}^{*}$.
The above theorem demonstrates the importance of maximum EFR: it exactly contains all states that can be rendered safe in real-time domain. We have a similar conclusion for EFR of a policy: if a state $x$ is in the EFR of a policy $\pi$, it can be rendered safe by this policy. This is formally stated in the following theorem, which theoretically justifies applying a policy in its EFR.
Theorem 4.2.6.
@@
Proof 4.2.7.
According to Definition 4.2.2,
@@
where $x_{t+i}$ are obtained by $\pi$. Since $\mathrm{X}_{\mathrm{init}}^{g}\subseteq\mathrm{X}_{\mathrm{cstr}}$, we have $x_{t+i}\in\mathrm{X}_{\mathrm{cstr}}$, i.e., $h(x_{t+i})\leq 0$.
Theorem 4.2.6 tells us that the EFR of a policy is where we can safely apply the policy in real-time domain. This is the reason why we need to find not only the policy but also its EFR when solving constrained OCPs. In literature, this EFR is found by identifying a safety certificate, e.g., control barrier function (CBF) [ames2014control], safety index (SI) [liu2014control], and Hamilton-Jacobi (HJ) reachability function [mitchell2005time]. When solving a constrained OCP, we aim to find a policy with a maximum EFR. This is because the size of EFR not only determines the policy’s working region but also affects its performance. If we want a policy to be optimal, its EFR must contain all optimal trajectories. A larger EFR has a greater chance of containing optimal trajectories and therefore increases the probability of finding the optimal policy.
However, it is important to note that existing methods still face significant challenges in finding the maximum EFR. For HJ reachability, the primary issue is scalability. While recent approaches using neural networks and RL offer promising paths to mitigate this [fisac2019bridging, yu2022reachability], they often lack formal guarantees. Methods that combine neural network verification with HJ analysis aim to provide these guarantees but often face scalability limitations again [yang2025scalable], making it an open problem. For CBFs and SIs, these approaches do not in general provide methods to compute the maximum EFR. Synthesizing a CBF or SI that certifies the maximum EFR typically requires guidance from other tools, such as HJ reachability, to shape or tune its parameters [yang2023synthesizing], since these methods inherently rely on an a priori characterization of the EFR.
Chapter 5 Some Properties of Feasible Regions
In this chapter, we formally explore the properties of feasible regions, as well as their relationships and how they are informative for virtual-time constraint design. Specifically, we will pay attention to containment relationships and equivalence relationships of different feasible regions.
@@ 5.1 Containment relationships of feasible regions
The definitions of feasible regions can be categorized by two dimensions: a) initial ones or endless ones and b) with a given policy or without a given policy. The four combinations correspond to 1) IFR $\mathrm{X}_{\mathrm{init}}^{g}$, 2) IFR of a policy $\mathrm{X}_{\mathrm{init}}^{g}(\pi)$, 3) EFR $\mathrm{X}_{\mathrm{edls}}^{g}$, and 4) EFR of a policy $\mathrm{X}_{\mathrm{edls}}^{g}(\pi)$. Besides these, we also need to pay attention to the constrained set $\mathrm{X}_{\mathrm{cstr}}$ and the maximum EFR $\mathrm{X}_{\mathrm{edls}}^{*}$. Among these six regions, $\mathrm{X}_{\mathrm{cstr}}$, $\mathrm{X}_{\mathrm{init}}^{g}$, and $\mathrm{X}_{\mathrm{init}}^{g}(\pi)$ are related only to the current step, while $\mathrm{X}_{\mathrm{edls}}^{g}$, $\mathrm{X}_{\mathrm{edls}}^{g}(\pi)$, and $\mathrm{X}_{\mathrm{edls}}^{*}$ are related to infinite future steps, and finding them are our ultimate goals. Since the latter three regions are related to long-term feasibility, their identification is much more difficult than that of the former. The following theorem gives some containment relationships between these six regions so that we can let those easy-to-identify regions cast some light on our goal regions.
Theorem 5.1.1 (Feasible region containment).
@@
    The first containment relationship $\mathrm{X}_{\mathrm{edls}}^{g}\subseteq\mathrm{X}_{\mathrm{init}}^{g}$ can be directly concluded from Definition 4.2.1. For the second one, according to Definition 4.1.1,
@@
    This holds by Definition 4.2.2 and Definition 4.1.2.
@@
    where $x_{t}=x$. Thus, $x_{t+i}\in\mathrm{X}_{\mathrm{init}}^{g}(\pi)$. Since $\mathrm{X}_{\mathrm{init}}^{g}(\pi)\subseteq\mathrm{X}_{\mathrm{init}}^{g}$, it follows that $x_{t+i}\in\mathrm{X}_{\mathrm{init}}^{g}$. Thus, $x_{t+i}\in\mathrm{X}_{\mathrm{cstr}}$, i.e., $h(x_{t+i})\leq 0$. Recall Theorem 4.2.4 and we reach $\mathrm{X}_{\mathrm{edls}}^{g}(\pi)\subseteq\mathrm{X}_{\mathrm{edls}}^{*}$.
@@
    Since $\mathrm{X}_{\mathrm{init}}^{g}=\mathrm{X}_{\mathrm{init}}^{g}(\pi)$, it holds that $\pi$ is initially feasible in all states in $\mathrm{X}_{\mathrm{init}}^{g}$. That is to say, $\pi\in\Pi_{\mathrm{init}}^{g}(x)$$\forall x\in\mathrm{X}_{\mathrm{init}}^{g}$. Starting from any $x\in\mathrm{X}_{\mathrm{edls}}^{g}$ and choosing
@@
Figure 5.1(a)-(d) illustrate the four containment relationships in Theorem 5.1.1 respectively. The light-colored circles correspond to virtual-time states, and the dark-colored ones stand for real-time states. The proof of the first two relationships can be easily understood by definition, while that of the last two needs some additional explanation. The left of Figure 5.1(c) shows the zero-sublevel sets of $h(x)$ and $g(x)$, which correspond to the maximum EFR $\mathrm{X}_{\mathrm{edls}}^{*}$ and the EFR of a policy $\mathrm{X}_{\mathrm{edls}}^{g}(\pi)$ on the right, respectively. To see how $\{x|h(x)\leq 0\}$ is related to $\mathrm{X}_{\mathrm{edls}}^{*}$, recall Theorem 4.2.4 which states that $x\in\mathrm{X}_{\mathrm{edls}}^{*}$ is equivalent to the existence of a policy rendering $x$ safe in real-time domain. Likewise, $x\in\mathrm{X}_{\mathrm{edls}}^{g}(\pi)$ means that policy $\pi$ renders $x$ safe in virtual-time domain, without violating virtual-time constraint $g(x)\leq 0$. Naturally, a larger valid state set ($\{x|h(x)\leq 0\}$ versus $\left\{x|g(x)\leq 0\right\}$) yields a larger EFR. The left of Figure 5.1(d) corresponds to the condition $\mathrm{X}_{\mathrm{init}}^{g}(\pi)=\mathrm{X}_{\mathrm{init}}^{g}$, with the red region representing both $\mathrm{X}_{\mathrm{init}}^{g}$ and $\mathrm{X}_{\mathrm{init}}^{g}(\pi)$. This condition implies that $\pi\in\Pi_{\mathrm{init}}^{g}(x),\forall x\in\mathrm{X}_{\mathrm{init}}^{g}$. The right picture illustrates how $\mathrm{X}_{\mathrm{edls}}^{g}$ is a subset of $\mathrm{X}_{\mathrm{edls}}^{g}(\pi)$. For a state in $\mathrm{X}_{\mathrm{edls}}^{g}$, every initially feasible policy at every time step must lead to an initially feasible state. As illustrated in the figure, policy $\pi^{\prime}$ (in purple) and $\pi^{\prime\prime}$ (in red) are two arbitrarily chosen policies, and $\pi$ is also a possible choice for $\pi^{\prime}$ and $\pi^{\prime\prime}$ (and all future policies). This guarantees that $\pi$ renders any state in $\mathrm{X}_{\mathrm{edls}}^{g}$ endlessly feasible.
Figure 5.1: Illustration of containment relationships.